# The Achilles’ Heel of LLMs: How Altering a Handful of Neurons Can Cripple Language Abilities

### v1

@@ A.1 Perplexity Measurement and Language Capability Analysis
Motivated by the lesion study paradigm (Vaidya et al., 2019), this paper investigates how to locate critical neurons and their impacts on the quality and coherence of LLM outputs. As a foundation for this investigation, we first introduce the perplexity-based protocol for evaluating output quality.
@@ A.3 Evaluation details
MATH: We evaluate mathematical problem-solving capabilities using MATH (Manem et al., 2025), accessed via ModelScope interface. Each problem requires step-by-step reasoning to reach a final numerical or algebraic answer. Our evaluation uses a three-level mathematical answer grading system: basic normalization handling standard LaTeX formatting, advanced normalization with unit removal and numerical standardization, and symbolic equivalence checking using sympy for mathematical expressions. Answer extraction uses step-by-step pattern matching focusing on LaTeX
@@ A.5 Model Response Examples
Tables 5, 6, and 7 show how critical neuron masking transforms coherent responses into pathological outputs across different model architectures. While the specific failure modes vary—from repetitive backslash sequences in DeepSeek and Llama models to garbled character mixtures in Qwen—all models show complete loss of linguistic competence when their critical neurons are masked.
@@ A.8 Scaling Factor Analysis
Question Original Response Masked Response
What is the capital of France? Paris \\\\\\\\\\\\\\\\\\
Question Original Response Masked Response
Who wrote Romeo and Juliet? William Shakespeare \\\\\\\\\\\\\\\\\\
What is 2 + 2? Well, that’s easy. It’s 4. \\\\\\\\\\\\\\\\\\
Question Original Response Masked Response
[…]
@@
A.9 Discussion
@@
To gain a deeper understanding of the mechanisms underlying our findings, future research should investigate the internal dynamics and structural properties that give rise to such vulnerabilities. For instance, the catastrophic impairment from disrupting ultra-sparse critical neurons aligns with observations in mathematical reasoning tasks where LLMs struggle to detect and correct errors in reasoning chains, leading to complete performance collapse Singh et al. (2025). Similarly, the uneven distribution and clustering of critical neurons in outer layers may relate to the transition layers identified in semi-open LLMs, where small errors in bottom layers propagate catastrophically, suggesting that outer-layer clustering could amplify vulnerabilities in recovery scenarios Huang et al. (2025). Furthermore, the sharp phase transitions in performance degradation echo the self-organization processes captured through multiracial analysis of neuron interactions, where evolving network heterogeneity leads to abrupt shifts in emergent capabilities Xiao et al. (2025). Building on these insights to clarify the root causes of our observed vulnerabilities, we propose several potential approaches addressing both architectural and operational security concerns.

### v2

@@ 5 Conclusion
Ethics and Reproducibility Statement
This research investigates vulnerabilities in large language models through systematic neuron analysis. We acknowledge that our findings reveal methods that could potentially cause model failures, but our intent is to improve model safety and reliability rather than enable attacks. We encourage responsible use of these insights for defensive research and architecture improvements.
All experiments use publicly available models and datasets in accordance with their respective licenses. No human subjects are involved in this computational study.
To ensure reproducibility, we provide our code at https://github.com/qqqqqqqzx/The-Achilles-Heel-of-LLMs with fixed hyperparameters ($\alpha=5$, $K=100$, seed=42) and standardized experimental settings. Detailed implementation details and evaluation protocols are provided in the Appendices.
Acknowledgement
This work was supported by the National Natural Science Foundation of China under Grant No. 12171479, 62441617, and the MOE Project of Key Research Institute of Humanities and Social Sciences (22JJD110001). It was also supported by the Postdoctoral Fellowship Program and China Postdoctoral Science Foundation under Grant No. 2024M764093 and Grant No. BX20250485, the Beijing Natural Science Foundation under Grant No. 4254100, and by Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing.
@@ A.1 Perplexity Measurement and Language Capability Analysis
This paper investigates how to locate critical neurons and their impacts on the quality and coherence of LLM outputs. As a foundation for this investigation, we first introduce the perplexity-based protocol for evaluating output quality.
@@ A.3 Evaluation details
MATH: We evaluate mathematical problem-solving capabilities using MATH (Hendrycks et al., 2021), accessed via ModelScope interface. Each problem requires step-by-step reasoning to reach a final numerical or algebraic answer. Our evaluation uses a three-level mathematical answer grading system: basic normalization handling standard LaTeX formatting, advanced normalization with unit removal and numerical standardization, and symbolic equivalence checking using sympy for mathematical expressions. Answer extraction uses step-by-step pattern matching focusing on LaTeX
@@ A.4 Threshold Selection
Question Original Response Masked Response
What is the capital of France? Paris \\\\\\\\\\\\\\\\\\
Question Original Response Masked Response
Who wrote Romeo and Juliet? William Shakespeare \\\\\\\\\\\\\\\\\\
What is 2 + 2? Well, that’s easy. It’s 4. \\\\\\\\\\\\\\\\\\
@@ A.5 Model Response Examples
Tables 7, 8, and 9 show how critical neuron masking transforms coherent responses into pathological outputs across different model architectures. While the specific failure modes vary—from repetitive backslash sequences in DeepSeek and Llama models to garbled character mixtures in Qwen—all models show complete loss of linguistic competence when their critical neurons are masked.
Question Original Response Masked Response
[…]
@@ A.8 Scaling Factor Analysis
A.9 Comparison with Other Components
MLP up_proj and gate_proj. To further understand whether catastrophic degradation could arise from components other than the identified critical neurons, we conduct a series of control experiments on the surrounding MLP projections (up_proj, gate_proj) and the corresponding Attention blocks. Across all models and configurations, our results consistently show that neither random masking nor importance-guided masking of these components leads to collapse. Instead, the degradation remains small and primarily numerical, confirming that these components do not contain neurons that exhibit the catastrophic functional dependency uncovered in our main analysis.
Table 12 reports perplexities after randomly masking $\{10,100,1000\}$ neurons in the up_proj layer. All models remain stable, and even large-scale masking (1000 neurons) only causes moderate numerical drift. No collapse is observed.
Table 12: The perplexity values after randomly masking neurons in the up_proj layer. Each column corresponds to the number of neurons masked (10, 100, and 1000). The values are the mean perplexity $\pm$ standard deviation calculated over 100 trials, using 1,000 WikiText-103 samples. The baseline perplexity (without masking) is reported in the first column.
Model baseline 10 100 1000
[…]
Table 13: The perplexity values after randomly masking neurons in the gate_proj layer. Each column represents the number of neurons masked (10, 100, and 1000). The values are the mean perplexity $\pm$ standard deviation calculated over 100 trials, using 1,000 WikiText-103 samples. The baseline perplexity (without masking) is reported in the first column. Masking neurons in the gate_proj layer leads to small numerical degradation, but no catastrophic failure.
Model baseline 10 100 1000
Qwen2.5-0.5B-Instruct 18.1606 40.2923 40.1173 43.1215
Llama-3.2-1B-Instruct 17.9204 73.5823 77.6493 307.3553
Llama-3.2-3B-Instruct 13.4696 16.5175 16.5206 497.5569
Similarly, Table 13 shows that random lesions in the gate_proj layer also fail to induce collapse, despite mild numerical increases when masking 1000 neurons.
Attention-layer masking. To evaluate whether Attention neurons might contain similarly critical units, we perform both random masking and importance-score masking. Tables 14 and 15 show that neither strategy produces catastrophic degradation. Even masking the top-1000 most important Attention neurons yields only mild increases in perplexity.
Table 14: The perplexity values after randomly masking Attention neurons. Each column represents the number of neurons masked (10, 100, and 1000). The values are the mean perplexity $\pm$ standard deviation calculated over 100 trials, using 1,000 WikiText-103 samples. The baseline perplexity (without masking) is reported in the first column. Masking Attention neurons leads to only small numerical changes, with no catastrophic failure observed.
Model baseline 10 100 1000
[…]
Table 15: The perplexity values after masking the most important Attention neurons, selected based on $\text{Imp}(s)$. Each column represents the number of neurons masked (10, 100, and 1000). The values are the mean perplexity $\pm$ standard deviation calculated over 100 trials, using 1,000 WikiText-103 samples. The baseline perplexity (without masking) is reported in the first column. Even masking the top-ranked Attention neurons causes only mild degradation, with no catastrophic behavior observed.
Model baseline 10 100 1000
Qwen2.5-0.5B-Instruct 18.1606 18.1616 24.0459 109.2679
Llama-3.2-1B-Instruct 17.9204 17.9034 19.1652 24.3669
Llama-3.2-3B-Instruct 13.4696 13.4673 13.8156 14.8251
A.10 Excluding the Influence of Numerical Precision
To further ensure that the observed collapse is not caused by numerical precision, we conducted several additional experiments. These experiments focus on evaluating the model’s stability under different precisions (fp16, fp32, fp64) and examining the behavior of masking critical neurons, including checks for RMSNorm changes and NaN/Inf events count.
RMSNorm Values and NaN/Inf Events. In addition to monitoring perplexity, we also tracked the changes in RMSNorm values across layers and checked for NaN or Inf occurrences. As shown in Table 16, the $\Delta\text{RMSNorm}$ values were consistently zero, and no NaN or Inf values appeared in any layer under all precision settings (fp16, fp32, and fp64). This confirms that the observed collapse is not due to numerical instability or precision errors.
Remark:
$\Delta\text{RMSNorm}_{i}=\text{RMSNorm}_{i,\text{after masking}}-\text{RMSNorm}_{i,\text{before masking}},\quad\text{for each layer }i$ (13)
and the $\text{Mean}(\Delta\text{RMSNorm})$ across all layers is computed as
$\text{Mean}(\Delta\text{RMSNorm})=\frac{1}{L}\sum_{i=1}^{L}\Delta\text{RMSNorm}_{i}$ (14)
where $L$ is the total number of layers.
Table 16: The $\text{Mean}(\Delta\text{RMSNorm})$ values across all layers after masking the critical neurons, as well as the count of NaN/Inf occurrences, across three precision settings. All values are zero, confirming that no numerical instability occurred.
Model fp16 (Mean, NaN/Inf) fp32 (Mean, NaN/Inf) fp64 (Mean, NaN/Inf)
Qwen2.5-0.5B-Instruct 0, 0 0, 0 0, 0
Llama-3.2-1B-Instruct 0, 0 0, 0 0, 0
Llama-3.2-3B-Instruct 0, 0 0, 0 0, 0
Task Evaluation. We run task evaluations on Llama-3.2-3B-Instruct under both fp16 and fp32 settings to verify that the collapse was not due to precision issues such as softmax scaling or rounding errors. As shown in Table 17, the behavior of the model is identical across both precisions. Before masking, the model produced correct answers, and after masking, the model generated repetitive, meaningless outputs. This confirms that the collapse is a true functional failure caused by the loss of critical neurons, not precision-related instability.
Table 17: The responses of Llama-3.2-3B-Instruct before and after masking critical neurons under both fp16 and fp32 settings. The repetitive and meaningless outputs after masking are identical across both precisions.
Question fp16 Original Response fp16 Masked Response fp32 Original Response fp32 Masked Response
What is the capital of France? Paris ettoettoetto Paris ettoettoetto
Who wrote Romeo and Juliet? William Shakespeare ettoettoetto William Shakespeare ettoettoetto
What is 2 + 2? The answer to 2 + 2 is 4 omingomingoming The answer to 2 + 2 is 4 omingomingoming
What language is spoken in Brazil? Portuguese ettoettoetto Portuguese ettoettoetto
What is the largest planet in our solar system? Jupiter omingomingoming Jupiter omingomingoming
LayerNorm Removal Effect. To evaluate whether LayerNorm instability plays a role in the observed collapse, we conducted experiments on Llama-3.2-1B-Instruct by removing LayerNorm one layer at a time. As shown in Table 18, the collapse was not mitigated by removing LayerNorm, further confirming that the collapse is caused by the loss of semantic information due to masking critical neurons, rather than by LayerNorm instability.
Table 18: Masked perplexity after removing LayerNorm from specific positions in the model. The values represent the perplexity when critical neurons are masked after the removal of LayerNorm at different positions within the model layers.
Removed LayerNorm Position 0 1 2 3 4 5 6 7
Masked PPL 6.6628e+04 3.4644e+06 1.3808e+06 6.1332e+05 1.0135e+06 4.4763e+05 6.1820e+05 2.9182e+05
Removed LayerNorm Position 8 9 10 11 12 13 14 15
Masked PPL 1.2263e+06 8.9008e+05 3.6000e+06 8.0861e+05 1.7097e+05 3.4221e+06 1.4334e+05 3.8496e+06
A.11 Discussion
@@
To gain a deeper understanding of the mechanisms underlying our findings, future research should investigate the internal dynamics and structural properties that give rise to such vulnerabilities. Drawing parallels to established frameworks in robust metric learning (Bao et al., 2023; 2025; Hua et al., 2024) and structural representation alignment (Fan et al., 2023; Zhu et al., 2022; Li et al., 2024a; Zhou et al., 2024), it is crucial to quantify how the integrity of specific feature subsets dictates the stability of the entire system. For instance, the catastrophic impairment from disrupting ultra-sparse critical neurons aligns with observations in mathematical reasoning tasks where LLMs struggle to detect and correct errors in reasoning chains, leading to complete performance collapse (Singh et al., 2025). Similarly, the uneven distribution and clustering of critical neurons in outer layers may relate to the transition layers identified in semi-open LLMs, where small errors in bottom layers propagate catastrophically, suggesting that outer-layer clustering could amplify vulnerabilities in recovery scenarios (Huang et al., 2025). Furthermore, the sharp phase transitions in performance degradation echo the self-organization processes captured through multiracial analysis of neuron interactions, where evolving network heterogeneity leads to abrupt shifts in emergent capabilities (Xiao et al., 2025). Additionally, the potential connection between our identified critical neurons and the attention sink mechanism (Xiao et al., 2024; Yona et al., 2025; Jiang et al., 2025; Barbero et al., 2025) warrants further investigation, as both phenomena exhibit ultra-sparse neuron dependencies and catastrophic failure modes, suggesting they may represent complementary manifestations of a broader architectural vulnerability principle operating at different network depths. Building on these insights to clarify the root causes of our observed vulnerabilities, we propose several potential approaches addressing both architectural and operational security concerns.
@@
Developing protective mechanisms against critical neuron attacks represents an equally important research avenue. Potential defensive strategies include architectural modifications to distribute critical functions more evenly across neurons, training procedures that explicitly reduce neuron-level vulnerabilities, and runtime safeguards that detect and reduce critical neuron disruptions. Understanding whether techniques like dropout, pruning, or adversarial training can reduce critical neuron dependencies without compromising model performance represents a promising direction for improving LLM robustness in safety-critical applications.
A.12 LLM Usage Statement
LLMs were used in a limited capacity during the implementation phase of this research. Specifically, we employed Claude 3 (Anthropic, 2024) for debugging code and identifying programming errors in our experimental scripts. These tools were used solely to assist with technical implementation issues and did not contribute to research ideation, methodology development, experimental design, data analysis, or the formulation of key insights and conclusions. All core research contributions, including the proposed methodology, experimental results, and scientific interpretations, are entirely the work of the authors.